OpenStudy (rebecca1588):

Which ordered pair is the solution to the system of equations? x+3y=12 -4x+3y=-3 A.(0, 4) B.(6, 2) C.(3, 3) D.(0, −1)

OpenStudy (tkhunny):

Have you considered substituting the given values into the two equations?

OpenStudy (rebecca1588):

I think its C.(3, 3)?

OpenStudy (welshfella):

are you guessing or did you work it out?

OpenStudy (tkhunny):

Demonstrate the substitutions and the results.

OpenStudy (rebecca1588):

I worked it out but I'd just like a second opinion :)

OpenStudy (tkhunny):

SHOW US! We can't know what you did if you don't show us. Give us the 1st Opinion and we can provide the second.

OpenStudy (rebecca1588):

-x-3y=-12 -4x+3y=-3 ----------- -5x=-15 x=3 and then 1 3+3y=12 3y=9 y=3

OpenStudy (rebecca1588):

I'm going with C Thanks anyways.

OpenStudy (welshfella):

good work

OpenStudy (tkhunny):

Excellent. Next time, start with that.